johns age is 13yrs greater than 3times dons age. ron is 5yrs younger than don. if the difference between ten times rons age and johns age is 28.
johns age?
rons age ?

Answer by mananth(16949)

johns age is 13yrs greater than 3times dons age. ron is 5yrs younger than don. if the difference between ten times rons age and johns age is 28.
johns age?
rons age ?
Don;s age x
John's age = 3x+13
Ron = x-5
..
10(x-5)=3x+13+28
10x-50=3x+13+28
10x-3x = 50+41
7x =91
x =91/7
x= 13
don's age
john's age - 3x+13
= 3*13+13
=52 years
...
Rons age = x-5
=13-5
=8 years